OpenStudy (anonymous):

Standard Form of the equation y = 1/3x - 4.

OpenStudy (ganpat):

that is in standard form.. y = -mx + c...

OpenStudy (anonymous):

Standard form is Ax + By + C = 0

OpenStudy (anonymous):

Slope intercept form... y = mx + b

OpenStudy (ganpat):

ha ha ha.. so u already have the answer.. LOL !!! compare it with ur equation :-p

OpenStudy (anonymous):

i dont know it

OpenStudy (anonymous):

y = 1/3 x - 4 <=== slope intercept form . . . multiply both sides by 3 to get rid of the fraction 3y = x - 12 . . . move x to the other side x - 3y = -12 <==== standard form

OpenStudy (anonymous):

thank you